Β Write the normalized Jones vectors for each of the following waves, and describe completely the state of polarization of each. (a) 𝐸 βƒ— = πΈπ‘œ cos(π‘˜π‘§ βˆ’ πœ”π‘‘)π‘₯ Μ‚ βˆ’ πΈπ‘œ cos(π‘˜π‘§ βˆ’ πœ”π‘‘)𝑦 Μ‚Β Β   
(b) 𝐸 βƒ— = πΈπ‘œ sin2Ο€(𝑧/lamda βˆ’ 𝑣𝑑)π‘₯ Μ‚ + πΈπ‘œ sin2Ο€( 𝑧/lamda βˆ’ 𝑣𝑑)𝑦 Μ‚Β 
1
Expert's answer
2021-04-26T17:13:25-0400

a)

"\vec{E}=E_x\vec{x}+E_y\vec{y},"

"E_x=-E_y," here phase difference between x- and y-components is "\pi,"

"E=\begin{pmatrix}
 E_x\\
 E_y
\end{pmatrix}=E_0\begin{pmatrix}
 1 \\
 -1
\end{pmatrix},"

"E^*\cdot E=1"

"E_0^2(1^2+(-1)^2)=1,"

"E_0=\frac{1}{\sqrt 2},"

normalized Jones vector will be

"\frac{1}{\sqrt2}\begin{pmatrix}
 1 \\
 -1
\end{pmatrix}," hence it represents linearly –45Β° polarized light.


b)

"\vec{E}=E_x\vec{x}+E_y\vec{y},"

"E_x=E_y," here phase difference between x- and y-components is "0,"

"E=\begin{pmatrix}
 E_x\\
 E_y
\end{pmatrix}=E_0\begin{pmatrix}
 1 \\
 1
\end{pmatrix},"

"E^*\cdot E=1"

"E_0^2(1^2+1^2)=1,"

"E_0=\frac{1}{\sqrt 2},"

normalized Jones vector will be

"\frac{1}{\sqrt2}\begin{pmatrix}
 1 \\
 1
\end{pmatrix}," hence it represents linearly 45Β° polarized light.